Suppose A is a square matrix such that det(A) = 2 and det(3A the power of t) = 18 then find the order of matrix A
1
Expert's answer
2020-01-30T08:39:20-0500

Suppose "A \in M_n(\mathbb R)."

Then, "\det (3A^t) = 3^n\det(A^t) = 3^n\det(A^t)="

"=3^n\det(A)=2 \cdot 3^ n=18,"

because the determinant is multilinear as a function of rows and the determinant respects the matrix multiplication.

We have, then,

"3^n = 9,"

which implies that "n=2" .

The solution relies on the assumption that "t \in \mathbb Z" because "A^t" might not be defined otherwise.